Please show work of Physical Chemistry: Estimate the value of P for the harpoon reaction between K and Cl2 for which I = 420 kJ/mol, Ea = 250 kJ/mol, radius atom K = 0,24 nm, and radius molecule Br2 = 228 pm, d = 400 pm.

Answer key : P = 4.2
Answer #1

Harpoon Reaction a type of Chemical Reaction between metal S & f hologen Reaction involves an strong electropositive of elect.

Harpoon Reaction a type of Chemical Reaction between metal & fhologen Reaction involves an strong electropositive of electronegative Element. O The Reaction is K + Br2 KBr + Bri P = o - po steric factor * = TR2 fra T d. - d = Rk + R Bx 2 je so plyte & CI-Ea)) e = 1.6810c . d = 400 x 10-12 T = 3.14 Eo : 8-8x 10-12 = I= 420 KI mot Ea = 250 Ilma P = (1.6810-19 4x3.14%818x10 12x400x16" (420-250)xit3 4.2